# MT-RewardTree: A Comprehensive Framework for Advancing LLM-Based Machine Translation via Reward Modeling

### v1

@@ 6.1 Test-time Alignment
Task setup. Test-time alignment, also known as decoding-time alignment (Huang et al., 2024), refers to the process of adjusting an LM’s output during inference to better align with human preferences, without additional training or fine-tuning. Its application in MT remains underexplored.
@@
Results. We use Qwen2.5-14B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct for generating tokens and leverage MT-PRM-LLaMA-3.2-3B and MT-PRM-Qwen-2.5-3B as the models for providing token-level rewards. We randomly sample 500 cases from the WMT 2023 testset. As shown in Figure 3, the reward-guided decoding methods outperform the standard greedy decoding in both EN-RU and ZH-EN translation tasks, evaluated by the COMET (Rei et al., 2020), COMETKiwi (Rei et al., 2022), and XCOMET-XL (Guerreiro et al., 2024) metrics. For instance, using the XCOMET-XL metric, LLaMA PRM and Qwen PRM outperform the standard greedy decoding by 17.5% and 17.9% in the EN-RU task respectively. Additionally, Qwen PRM slightly outperforms LLaMA PRM in both translation tasks and across all metrics, which aligns with the results in Table 3.1, where Qwen PRM achieves better token-level reward performance. These findings highlight the effectiveness of reward-guided decoding strategies in improving MT outcomes.
Case Study. The case in Table 5 compares a ZH-EN translation using Greedy Decoding (GD), GD with LLaMA PRM, and GD with Qwen PRM. The standard GD translation, “The mechanic said the oil was clear, indeed good oil” , conveying the basic meaning but lacking the important "changing oil" context. Both reward-guided decoding methods improve the translation to a more complete, high-quality version, achieving a COMETKiwi score improvement of 4.97%.
@@
6.2 Hypothesis Ensembling
Task Setup. Ensembling is widely recognized for its ability to combine multiple complementary models to improve performance in machine learning (Hansen and Salamon, 1990; Ting and Witten, 1997). Farinhas et al. (2023) explore generating multiple translation hypotheses and ensembling them to produce a single higher-quality translation. In this work, we investigate two complementary ensembling approaches for MT: ranking-based selection and Minimum Bayes Risk decoding.
In ranking-based ensembling, the optimal output is selected via:
$\hat{y}_{\mathrm{ranking}}:=\mathbf{argmax}_{y\in\mathcal{\bar{Y}}}~{}f(y)$ (14)
where $f$ represents either: (1) Strong LLM as a judge (e.g., Gemini-2.0-Flash [^6] https://ai.google.dev/gemini-api/docs/models) or (2) Our PRM, converted to sequence-level scoring through weighted DPO rewards (as shown in Eq. 11).
For MBR decoding, we maximize the expected utility over candidate references:
$\hat{y}_{\mathrm{mbr}}:=\mathbf{argmax}_{y\in\bar{\mathcal{Y}}}\frac{1}{M}\sum_{j=1}^{M}u(y^{(j)},y)$ (15)
where $u(\cdot)$ can be instantiated using traditional metrics such as BLEU (Papineni et al., 2002), BERTScore (Zhang et al., 2019), and the reference-based metric COMET (Rei et al., 2020) for upper-bound analysis. This dual-strategy framework allows for a comprehensive evaluation of the ensembling potential of reward models.
Results. We evaluate on 500 cases sampled from the WMT 2023 dataset, using the TowerInstruct-7B-v0.2 model with nucleus sampling to generate 8 candidate translations for each case. As shown in Table 5, our MT-PRM-LLaMA-3.2-3B outperforms MBR decoding, based on BLEU and BERTScore, by 0.93% and 0.27%, respectively. It even surpasses the commercial LLM Gemini-2.0-Flash by 0.51%. When comparing the performance of different PRMs, we find that LLaMA outperforms Qwen, which is consistent with their relative performance in the MT-RewardBench sequence-level evaluation for the Arbitrary set. This further validates the practical utility of MT-RewardBench.
@@ 7 Related Work
Token-Level Feedback Mechanisms. Fine-grained feedback has been recognized for its ability to help models capture potential errors more precisely (Lightman et al., 2024). In the context of mathematical reasoning, process supervision using Monte Carlo methods has shown significant promise (Wang et al., 2024c; Qi et al., 2024; Guan et al., 2025). Furthermore, developments in general-domain have demonstrated that DPO can implicitly learn token-level rewards through policy optimization, a process referred to implicit reward learning (Rafailov et al., 2024a; Wang et al., 2024a; Yuan et al., 2024). Despite these advancements, these approaches have yet to be tested in the context of MT. The translation community has long acknowledged the value of granular feedback, with early attempts relying on binary error markings from human annotations (Kreutzer et al., 2020), reference-based heuristics (Petrushkov et al., 2018), or LLM (Feng et al., 2024b).
Alignment Paradigms in Machine Translation. Alignment techniques in neural machine translation have evolved from Minimum Risk Training (Shen et al., 2015) to more sophisticated reinforcement learning approaches (Dang et al., 2024). While PPO-based RLHF has achieved success in general-domain alignment, its application to MT presents unique challenges, particularly due to the need for fine-grained quality signals rather than the bandit reward. Recent works like He et al. (2024) and Xu et al. (2024) have investigated the use of automatic metrics to select better translations or construct preference pairs to improve the LLM, while Zhao et al. (2024) explored scaling test-time compute to further enhance translation performance. Recently, Ramos et al. (2024) pioneered the use of xCOMET as a dynamic reward signal during RL training. However, these methods remain limited to sequence-level guidance or binary approximations of the reward process, failing to provide the fine-grained token-level feedback required for more accurate translation alignment. MT-RewardTree provides a new perspective by introducing a more granular, token-level reward modeling framework.
@@ 8 Conclusion
In this work, we propose MT-RewardTree, a comprehensive framework for constructing, evaluating, and deploying process reward models in machine translation. Our framework leverages an automatic token-level preference pair generation approach inspired by approximate Monte Carlo Tree Search, effectively addressing the challenge of large-scale fine-grained supervision annotation. Extensive experiments on both sequence-level and token-level benchmarks demonstrate that our MT-PRM achieves advanced performance in reward modeling in MT, surpassing traditional sequence-level preference pairs. Our exploration of test-time alignment and hypothesis ensembling provide valuable insights for the application of reward models.

### v2

@@ 5.3 Test-time Alignment
Test-time alignment, also known as decoding-time alignment (Huang et al., 2024; Rashid et al., 2024), refers to the process of adjusting an LM’s output during inference to better align with human preferences, without additional training or fine-tuning. Its application in MT remains underexplored.
@@
We use Qwen2.5-14B-Instruct [^3] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct for generating tokens and leverage MT-PRM-LLaMA-3.2-3B and MT-PRM-Qwen-2.5-3B as the models for providing token-level rewards. We randomly sample 500 cases from the WMT 2023 testset. As shown in Figure 3, the reward-guided decoding methods outperform the standard greedy decoding in both EN-RU and ZH-EN translation tasks, evaluated by the COMET (Rei et al., 2020), COMETKiwi (Rei et al., 2022), and XCOMET-XL (Guerreiro et al., 2024) metrics. For instance, using the XCOMET-XL metric, LLaMA PRM and Qwen PRM outperform the standard greedy decoding by 17.5% and 17.9% in the EN-RU task respectively. Additionally, Qwen PRM slightly outperforms LLaMA PRM in both translation tasks and across all metrics, which aligns with the results in Table 3.1, where Qwen PRM achieves better token-level reward performance. These findings highlight the effectiveness of reward-guided decoding strategies in improving MT outcomes.
A crucial aspect of reward-guided alignment is the trade-off with output diversity. Our framework explicitly manages this balance via the weighting parameter $w$ in the reward-guided scoring function (Eq. 13). The parameter $w$ controls the influence of the PRM: a higher $w$ prioritizes the learned preferences for stronger alignment, while a lower $w$ preserves more of the base model’s original output distribution, thereby maintaining diversity.
Model Task w=0 w=0.3 w=0.5 w=0.7
MT-PRM-Llama Guided ZH-EN 0.7167 0.8759 0.8420 0.8677
EN-RU 0.5984 0.7459 0.7619 0.7143
MT-PRM-Qwen Guided ZH-EN 0.7167 0.8723 0.8448 0.8671
EN-RU 0.5984 0.7444 0.7569 0.6721
Table 6: XCOMET scores showing the impact of the weighting parameter $w$ on test-time alignment. The baseline is Qwen2.5-14B-Instruct ($w=0$). Optimal values vary, allowing for tunable performance.
The quantitative results in Table 6 show that a range of $w$ values provides a significant boost over the baseline ($w=0$), allowing users to choose a balance that suits their needs. For example, with the Llama-PRM on EN-RU, $w=0.5$ yields the best XCOMET score, while $w=0.3$ is optimal for the MT-PRM-Qwen. Furthermore, the qualitative case study in Table 7 illustrates that even under PRM guidance, different models and weights produce varied, high-quality outputs, demonstrating that diversity can be maintained.
System Translation
Source (Reference) 放入充电盒中,耳机自动关机。 (Put it in the charging case, the headphones will turn off automatically.)
Qwen2.5-14B-Instruct ($w=0$) Put them in the charging case, and the headphones will shut off automatically.
MT-PRM-Llama ($w=0.3$) Put them in the charging case, and the headphones will shut off automatically.
MT-PRM-Qwen ($w=0.5$) When placed in the charging case, the headphones shut off automatically.
MT-PRM-Llama ($w=0.7$) Place the earbuds in the charging case, and they will shut down automatically.
Table 7: Qualitative case study illustrating output diversity under PRM guidance.
@@ 6 Related Work
Token-Level Feedback Mechanisms. Fine-grained feedback has been recognized for its ability to help models capture potential errors more precisely (Lightman et al., 2024). In the context of mathematical reasoning, process supervision using Monte Carlo methods has shown significant promise (Wang et al., 2024b; Qi et al., 2024; Guan et al., 2025). Furthermore, developments in general-domain have demonstrated that DPO can implicitly learn token-level rewards through policy optimization, a process referred to implicit reward learning (Rafailov et al., 2024a; Wang et al., 2024a; Yuan et al., 2024). Despite these advancements, these approaches have yet to be tested in the context of MT. The translation community has long acknowledged the value of granular feedback, with early attempts relying on binary error markings from human annotations (Kreutzer et al., 2020), reference-based heuristics (Petrushkov et al., 2018), or LLM (Feng et al., 2024b).
Alignment Paradigms in Machine Translation. Alignment techniques in neural machine translation have evolved from Minimum Risk Training (Shen et al., 2015) to more sophisticated reinforcement learning approaches (Dang et al., 2024). While PPO-based RLHF has achieved success in general-domain alignment, its application to MT presents unique challenges, particularly due to the need for fine-grained quality signals rather than the bandit reward. Recent works like He et al. (2024) and Xu et al. (2024) have investigated the use of automatic metrics to select better translations or construct preference pairs to improve the LLM, while Zhao et al. (2024) explored scaling test-time compute to further enhance translation performance. Recently, Ramos et al. (2024) pioneered the use of xCOMET as a dynamic reward signal during RL training. However, these methods remain limited to sequence-level guidance or binary approximations of the reward process, failing to provide the fine-grained token-level feedback required for more accurate translation alignment.
@@ 7 Conclusion
In this work, we propose MT-RewardTree, a comprehensive framework for constructing, evaluating, and deploying process reward models in MT. Our framework leverages an automatic token-level preference pair generation approach inspired by approximate Monte Carlo Tree Search, effectively addressing the challenge of large-scale fine-grained supervision annotation. Extensive experiments on both sequence-level and token-level benchmarks demonstrate that our MT-PRM achieves advanced performance in reward modeling in MT, surpassing traditional sequence-level preference pairs. Our exploration of token credit assignment and test-time alignment provide valuable insights for the application of reward models in MT.
@@
Acknowledgments
This work is supported by the National Key R&D Program of China (Grant No. 2024YFC3308304), the "Pioneer" and "Leading Goose" R&D Program of Zhejiang (Grant no. 2025C01128), the National Natural Science Foundation of China (Grant No. 62476241), the Natural Science Foundation of Zhejiang Province, China (Grant No. LZ23F020008), the Zhejiang Zhejiang Key Laboratory of Medical Imaging Artificial Intelligence.